# Kuan-Cheng Chen QuEST Imperial College London London, United Kingdom

### v2

@@ III-D2 Simulation of QSVM using cuStateVector
  • Algorithm 2 Get Kernel Matrix using statevector simulator
    Furthermore, we implement the simulation using the NVIDIA cuQuantum framework, replacing Algorithm 1 with calls to the cuStateVec library to accelerate state vector simulation on GPUs, described in Algorithm 2. The advantage of using cuStateVec includes a speedup of the simulation time by leveraging GPU capabilities and enabling multi-GPU processing with MPI for distributed computing. The effectiveness of cuStateVec in enhancing quantum-circuit-simulation efficiency is evidenced in Lykov et al.’s research work using cuStateVec and the cuQuantum SDK[22].
@@
      Calculate and store $kernel\_matrix[i_{1}-1][i_{2}-1]\leftarrow\sqrt{\text{amp.real}^{2}+\text{amp.imag}^{2}}$.
@@
  • Algorithm 3 Get Kernel Matrix using cuTensorNet with Network Context
@@
    III-D3 Simulation of QSVM using cuTensorNet
    However, even with opt-einsum facilitating CPU acceleration and cuStateVector enabling GPU acceleration, challenges persist due to the complexity of encoding the number of qubits $O(2^{n})$ and the size of the data $O(n^{2})$. To surmount these challenges, we present an innovative approach using the cuTensorNet library for QSVM simulation. In the creation of the tensor network representation, we seamlessly integrate Qiskit and cuQuantum’s built-in CircuitToEinsum object. Initially, Qiskit is used to construct a QuantumKernel circuit, which is then transformed into ’expression’ and ’operand’ components by CircuitToEinsum. Due to the identical topological structure of the quantum circuit, the same ’expression’ component can be reused for subsequent pairs of data. Meanwhile, the ’operand’ is updated with parameters from the previously created operand. This approach rapidly transitions data pairs into tensor networks and preserves computational efficiency. The derivation of the kernel matrix—a pivotal component of the SVM—exploits a consistent ’path’ to greatly minimize the repetition of contraction order calculations. The detailed algorithm is described in Algorithm 3. This technique not only leverages the computational strength of GPUs but also ensures path reusability, resulting in a considerable acceleration of the simulation process and a dramatic reduction in computational complexity. We will demonstrate those improvements in the next section.
    IV Performance and Benchmarking of QSVM with cuTensorNet
    IV-A QSVM Simulation and cuTensorNet-Accelerated QSVM (cuTN-QSVM)
    In the outlined simulation workflow, Fig. 1 and 5 illustrate the sequence from the initial input of data to the generation of a quantum circuit for the purpose of encoding. Subsequent steps involve the use of optimized compilation to compute and simulate the quantum circuits, leading to the extraction of a quantum kernel matrix. This matrix is then applied to develop a support vector classifier (SVC).
@@ IV-A QSVM Simulation and cuTensorNet-Accelerated QSVM (cuTN-QSVM)
    However, in typical CPU-based workflows, bottlenecks arise in the progression from the construction of quantum circuits to the calculation of the quantum kernel matrix, where the complexity of simulating the QSVM algorithm scales exponentially with the number of qubits, $O(2^{n})$, and quadratically with data size, $O(n^{2})$. To alleviate these bottlenecks, we incorporate the cuQuantum SDK into QSVM workflow, employing a method of assigned parameters for the formulation of QSVM’s quantum circuits. We then maintain a consistent ’expression’ for the simulation of these circuits. Ultimately, we apply a ’path reuse’ strategy within the tensor network to compute the quantum kernel matrix. This suite of strategies significantly mitigates the computational complexity associated with processing large datasets, reducing it from $O(n^{2})$ to $O(1)$ for pathfinding and substantially enhancing throughput over traditional CPU computations. Importantly, as depicted in Fig. 7, the expressions and paths used in the cuTensorNet during the QSVM simulation process remain unchanged compared to those in CPU and cuStateVector, ensuring that no accuracy is compromised for the sake of expedience. In addition to the path reuse strategy, cuTensorNet offers concurrent execution for tensor network contraction. This technique allows the continued contractions on the GPU asynchronously when tensors are already on the device, thus enhancing computational efficiency by continuing operations without delay. The pronounced speedup achieved through the implementation of path reuse within the cuTensorNet framework is detailed in Fig.7(g), where we report a fiftyfold increase in speed compared to conditions without path reuse.
@@
    In the comprehensive workflow outlined in Fig. 8, the input data initiates the construction of a quantum circuit, seamlessly integrated with Qiskit/Cirq. The process advances by converting quantum circuits into tensor networks represented as CuPy arrays, enabling the utilization of in-place operations to efficiently update content for the same operands. Key to enhancing computational efficiency within this framework is the strategic deployment of direct conversion from data to operand, alongside expression reuse for optimizing computational pathways. This step is crucial in minimizing redundancy and ensuring the streamlined execution of the workflow. As the process proceeds, CuPy’s capabilities are harnessed to accelerate the computation of the kernel matrix, culminating in the application of the SVC. Moreover, cuTensorNet, as part of the cuQuantum SDK, incorporates advanced strategies such as path reuse and non-blocking operations across multi-GPU configurations.
    These approaches significantly reduce the computational overhead from a conventional complexity of $O(2^{n})$ to a more scalable $O(n^{2})$, thereby enhancing the practicality of executing extensive QSVM simulations with improved processing times and efficiency in resource utilization. Fig. 9 illustrates that quantum simulation on the NVIDIA A100 GPU using cuStateVector becomes practically infeasible for more than 50 qubits. However, by employing cuTensorNet, single-contraction simulations can be completed within 0.2 seconds, even with up to 784 qubits. Additionally, Fig. 9 shows that the path reuse strategy can further enhance the speed, offering more than tenfold acceleration when increasing the number of qubits in the QSVM algorithm.
@@
    IV-B Accuracy Benchmarking
    Figure 10: Train and test accuracy of quantum and classical models as a function of the number of qubits. The plot compares quantum model performance with varying training set sizes (50 and 100 data points) against the optimal classical model.
    In our study’s accuracy benchmarking, we observed that the trained QSVM demonstrates an improvement in accuracy with increasing dataset sizes shown in Fig. 10. Additionally, there tends to be a positive correlation between the number of qubits and accuracy; a larger quantity of qubits typically yields higher accuracy. In the specific case under consideration, the optimized configuration entailed an algorithm using 128 qubits on a sample set of 100. This experiment also indicates that the accuracy sharply declines beyond roughly 200 qubits, attributed to the off-diagonal kernel matrix elements approaching zero. Under optimized parameters, the QSVM achieved test and training accuracies of 94% and 100%, respectively, outperforming the traditional SVM, which achieved accuracies of 90% in testing and 95% in training with a comparable amount of data. These empirical results underscore the enhanced capability of QSVM over traditional SVMs as data volume increases, especially in big data analytics applications, a finding that is corroborated by the work of Chen et al. [10].
    Furthermore, in Fig. 11, we extend our evaluation to larger data sizes, focusing on the MNIST dataset, particularly on digits 0 and 9, which are inherently challenging to distinguish. Our analysis reveals that with a limited training dataset size (i.e., 10 instances), the accuracy hovers near 50%. However, as the volume of training data surpasses 100 instances, the accuracy significantly increases to approximately 95%, surpassing the performance of the optimal classical SVM.
    Figure 11: The relationship between dataset size and accuracy for a 128-qubit QSVM algorithm, compared to the optimal classical model.
    IV-C Simulation with Single CPU and GPU
    In this section, we compare the performance of a CPU and a GPU, as illustrated in Fig.12. To ensure a fair comparison, we employed Opt-Einsum for the contraction process on a single AMD EPYC 7J13 CPU, contrasting this with a single NVIDIA A100 GPU using cuTensorNet for the contraction process, with path reuse implemented. The detailed pseudocodes are discussed in Section III-D. Moreover, it was necessary to synchronize the contraction paths in Opt-Einsum with those of cuTensorNet to ensure consistency. As depicted in Fig.12, the speedup provided by the GPU relative to the CPU becomes more pronounced as the number of simulated qubits increases. Consequently, for large-scale qubit simulations, GPUs demonstrate enhanced scalability and promise substantial benefits for future advanced qubit algorithms in simulation and emulation.
@@ IV-C Simulation with Single CPU and GPU
    Figure 12: Benchmark QSVM circuit simulation time using a single CPU and a single GPU.
    V Distributed simulation and Resource Estimation in HPC
    In the final section of our study, a multi-GPU instance was utilized to expand the QSVM model to accommodate a dataset comprising 1,000 data points of 28x28 MNIST images. The implementation of multi-GPU resources to enhance quantum circuit simulation via cuStateVector is thoroughly detailed in the research conducted by Shaydulin et al. [34]. Our emphasis lies on leveraging the data from these experiments to rigorously assess both the computational costs and the temporal demands inherent in simulating the QSVM algorithm within a multi-GPU processing framework.
    In our computational environment, each GPU within a node is interconnected using the high-bandwidth NVLink network, optimizing intra-node communication. Inter-node data transfer mandates that information from the GPU be first relayed to the CPU, then onward to the target node, a process that requires careful selection of a communication protocol sensitive to the GPU’s positional context. The MPI offers built-in capabilities for such operations, which are enabled by setting the $MPI\_GPU\_SUPPORT\_ENABLED$ environment variable. By harnessing these integrated technological benefits, we have successfully actualized the accelerated computational outcomes for managing large-scale qubit systems and extensive datasets, as illustrated in Fig.13. Comparative analysis indicates that our performance metrics are on par with distributed simulation results documented in the existing scientific corpus, as cited in Bayraktar et. al.’s and Lykov et. al.’s work[23, 22].
    V-A Benchmarking cuQuantum Multi-GPU with MPI
    Fig. 13 illustrates the execution time required for quantum simulations in relation to the number of qubits. The data compares the performance of a single A100 GPU to systems utilizing 2, 4 and 8 GPUs in conjunction with MPI and within a single NVIDIA DGX node. It is evident from the results that the incorporation of multi-GPUs significantly decreases computation time, highlighting the strong linear speedup of cuTenserNet with MPI. The trend indicates a substantial reduction in execution time as the number of GPUs is increased, affirming the efficacy of multi-GPU setups in handling large datasets.
@@ V-A Benchmarking cuQuantum Multi-GPU with MPI
    V-B Large Dataset Processing with Multi-GPU instance
@@
    VI Conclusion
    This paper has presented a comprehensive study on the application and efficacy of cuTN-QSVMs in conjunction with NVIDIA’s cuQuantum SDK. Our findings have consistently shown that by integrating cuTensorNet within our computational workflow, we have been able to significantly decrease the computational complexity involved in simulating QSVM algorithms, especially when working with large qubit counts. Through rigorous performance benchmarking, our approach has demonstrated not only enhanced simulation speeds and efficiency but also scalability across multi-GPU platforms. Additionally, we have observed substantial improvements in accuracy with increasing data set sizes, underlining the potential of quantum computing methods in tackling complex, large-scale data challenges. The integration of cuTensorNet and MPI-based multi-GPU systems reveals a path forward for quantum computing that offers practical advantages for both current research and future applications, bridging the gap between quantum hardware and classical computing resources. This points towards a promising direction for high-performance quantum simulations within the Quantum-HPC ecosystem. Furthermore, we utilized cuTN-QSVM to classify bone cell images captured by synchrotron transmission X-ray microscopy (TXM) at the BL01B1 beamline of the Taiwan Light Source (TLS) in Hsinchu, Taiwan. Bone cell image data possess a high-dimensional feature space but limited data quantity. Traditional machine learning struggles with classifying such data, but QSVM overcomes this, showcasing its advantages. Due to page limit, the classification results of bone cells will be elaborated in the conference.

### v3

@@ 3.4.1 Simulation of QSVM with state vector
  • Algorithm 1 Get Kernel Matrix using cuStateVec
    3.4.2 Simulation of QSVM with tensor network
    Even with cuStateVec enabling GPU acceleration, challenges persist due to the complexity of encoding the number of qubits $O(2^{n})$ and the size of the data $O(n^{2})$. To surmount these challenges, we present an innovative approach using the cuTensorNet library for QSVM simulation. In the creation of the tensor network representation, we seamlessly integrate Qiskit and cuQuantum’s built-in CircuitToEinsum object.
    Initially, Qiskit is used to construct a QuantumKernel circuit, which is then transformed into “expression” and “operand” components by CircuitToEinsum. Due to the identical topological structure of the quantum circuit, the same “expression” component can be reused for subsequent pairs of data. Meanwhile, the “operand” is updated with parameters from the previously created operand. This approach rapidly transitions data pairs into tensor networks and preserves computational efficiency. The derivation of the kernel matrix—a pivotal component of the SVM—exploits a consistent “path” to greatly minimize the repetition of contraction order calculations. The detailed algorithm is described in Algorithm 2. This technique not only leverages the computational strength of GPUs but also ensures path reusability, resulting in a considerable acceleration of the simulation process and a dramatic reduction in computational complexity. We will demonstrate those improvements in the next section.
@@ 3.4.2 Simulation of QSVM with tensor network
      Calculate and store
           $kernel\_matrix[i_{1}-1][i_{2}-1]\leftarrow\sqrt{\text{amp.real}^{2}+\text{amp.imag}^{2}}$.
@@
    Symmetrize $kernel\_matrix$ by adding its transpose and an identity matrix:
  • $kernel\_matrix\leftarrow kernel\_matrix+kernel\_matrix^{T}+\text{diag}(\mathbb{1}_{datasize1})$. return $kernel\_matrix$
    Algorithm 2 Get Kernel Matrix using cuTensorNet
    To ensure a fair comparison tensor network QSVM simulation between CPU and GPU performance, we utilize the opt-einsum package, which provides optimized tensor computation on CPUs similar to the cuQuantum SDK available for NVIDIA GPUs. The detailed algorithm for simulating the QSVM on CPUs, aimed at equalizing the computational environment to the extent possible, is described in Algorithm 3.
    1. 1.
    2. 2.
      Set the current operand index $i$ to $-1$.
    3. 3.
      for $i_{1},i_{2}\in\{1,\ldots,indices\}$ do
    1. (a)
      Update the operands index $i\leftarrow i+1$.
    2. (b)
      Compute amplitude
    $amp\leftarrow\text{opt\_einsum.contract}(exp,operands[i],path)$.
  • (c)
    Calculate and store
  •        $kernel\_matrix[i_{1}-1][i_{2}-1]\leftarrow\sqrt{\text{amp.real}^{2}+\text{amp.imag}^{2}}$.
    end for
  • 4.
@@
  • Algorithm 3 Get Kernel Matrix using opt-einsum
    4 Performance and Benchmarking of QSVM with cuTensorNet
    4.1 QSVM Simulation and cuTensorNet-Accelerated QSVM (cuTN-QSVM)
    In the outlined simulation workflow, Fig. 1 and 4 illustrate the sequence from the initial input of data to the generation of a quantum circuit for the purpose of encoding. Subsequent steps involve the use of optimized compilation to compute and simulate the quantum circuits, leading to the extraction of a quantum kernel matrix. This matrix is then applied to develop a support vector classifier (SVC).
@@ 4.1 QSVM Simulation and cuTensorNet-Accelerated QSVM (cuTN-QSVM)
    However, in typical CPU-based workflows, bottlenecks arise in the progression from the construction of quantum circuits to the calculation of the quantum kernel matrix, where the complexity of simulating the QSVM algorithm scales exponentially with the number of qubits, $O(2^{n})$, and quadratically with data size, $O(N^{2})$. To alleviate these bottlenecks, we incorporate the cuQuantum SDK into the QSVM workflow, employing a method of assigned parameters for the formulation of QSVM’s quantum circuits. We then maintain a consistent “expression” for the simulation of these circuits. Ultimately, a “path reuse” strategy is adopted for the tensor network contractions to compute the quantum kernel matrix, effectively reducing redundant computations when processing large datasets, reducing it from $O(N^{2})$ to $O(1)$ for pathfinding. Importantly, as depicted in Fig. 6, the expressions and paths used in the cuTensorNet during the QSVM simulation process remain unchanged compared to those in CPU and cuStateVec, ensuring that no accuracy is compromised for the sake of expedience. In addition to the path reuse strategy, cuTensorNet offers concurrent execution for tensor network contractions. This technique allows the continued contractions on multiple GPUs asynchronously when tensors are already on the device, thus enhancing computational efficiency by continuing operations without delay. The pronounced speedup achieved through the implementation of path reuse within the cuTensorNet library is detailed in Fig.6(g), where we report a fiftyfold increase in speed compared to conditions without path reuse.
    In the comprehensive workflow outlined in Fig. 7, the input data initiates quantum circuit construction, integrating frontends such as Qiskit or Cirq with the cuQuantum API, which generates Einstein summation expressions and tensor operands for the circuit. The process advances by converting quantum circuits into tensor networks represented as CuPy arrays, enabling the utilization of in-place operations to update content for the same operands efficiently. Key to enhancing computational efficiency within this framework is the strategic deployment of direct conversion from data to operand, alongside expression reuse for optimizing computational pathways. This step is crucial in minimizing redundancy and ensuring the streamlined execution of the workflow. As the process proceeds, CuPy’s capabilities are harnessed to accelerate the computation of the kernel matrix, culminating in the application of the SVC. Moreover, cuTensorNet, as part of the cuQuantum SDK, incorporates advanced strategies such as path reuse and non-blocking operations across multi-GPU configurations.
@@
    These approaches significantly reduce the computational overhead from a conventional complexity of $O(2^{n})$ to a more scalable $O(n^{2})$, thereby enhancing the practicality of executing extensive QSVM simulations with improved processing times and efficiency in resource utilization. Fig. 8 illustrates that quantum simulation on the NVIDIA A100 GPU using cuStateVec becomes practically infeasible for more than 50 qubits. However, by employing cuTensorNet, single-contraction simulations can be completed within 0.2 seconds, even with up to 784 qubits. Additionally, Fig. 8 shows that the path reuse strategy can further enhance the speed, offering more than tenfold acceleration when increasing the number of qubits in the QSVM algorithm.
@@
    4.2 Accuracy Benchmarking and Validation of Large-scale QSVM
    4.2.1 Binary Classification
    We benchmarked the performance of a QSVM on high-dimensional image classification tasks using both 10-class MNIST and Fashion-MNIST datasets of 31,500 images each, split 80–20 for training and testing. As a classical baseline, we employed an SVM with a radial basis function (RBF) kernel and systematically varied the scaling parameter $\gamma$ between 0.001 and 1000 to identify optimal hyperparameters. Although SVM is not the most advanced machine learning model available, it provides a well-established framework that enables us to validate the feasibility and potential advantages of our quantum-enhanced approach.
    Figure 9: Benchmark of QSVM and SVM on MNIST and Fashion-MNIST (nine labels, 45 binary classification tasks). (a) Training accuracy vs number of qubits (QSVM) or features (SVM). (b) Test accuracy. QSVM outperforms SVM for moderate qubit counts, but overfitting leads to reduced test accuracy with more qubits. Shaded regions show one standard deviation around mean accuracy.
    In Fig. 9, our results indicate that QSVM offers competitive performance and can, under certain conditions, outperform the classical SVM for moderate circuit sizes. In particular, QSVM maintains high accuracy by leveraging quantum feature maps that embed data into larger Hilbert spaces. However, beyond a critical qubit threshold, we observe a decline in test accuracy, which we attribute to overfitting effects and the phenomenon of Barren plateaus—where off-diagonal kernel matrix elements diminish and the optimization landscape becomes exponentially flat[40]. This vanishing-gradient problem not only complicates the training of parameterized quantum circuits but also underscores the practical limitations of naively scaling up circuit depth or qubit count.
    Figure 10: Classification Accuracy versus Dataset Size for Binary Classification of MNIST Digits 2 and 6 using a 128-Qubit QSVM model.
    In light of these observations, current quantum machine learning approaches still require careful feature engineering or hybrid methods to optimize model performance. Moreover, the amount of training data can significantly impact QSVM accuracy, as illustrated in Fig. 10, underscoring the importance of sufficiently large datasets.
    Despite these challenges, our proof-of-concept study confirms that QSVM can serve as a promising foundation for large-scale quantum machine learning, particularly in scenarios where high-dimensional embeddings may confer a computational advantage. Further research on quantum circuit design, regularization strategies, and optimization techniques will be crucial to fully harness the benefits of quantum-enhanced models and to mitigate the pitfalls associated with increasingly large quantum systems.
    4.2.2 Multi-class Classification
    Figure 11: Confusion matrices for the classical SVM with 64 features (a) and 96 features (c), and for the QSVM with 64 qubits (b) and 96 qubits (d), evaluated on 10-class MNIST (a, b) and 10-class Fashion-MNIST (c, d). Each cell indicates the number of predictions for a given true label (vertical axis) and predicted label (horizontal axis). Brighter diagonal entries reflect a higher count of correctly classified samples. Overall, QSVM demonstrates consistently strong performance, often improving upon the classical SVM.
    To assess the robustness of the QSVM beyond binary classification, we further evaluated its performance on the 10-class versions of the MNIST and Fashion-MNIST datasets. For each dataset, we select 1,000 training samples and reserve an additional 500 samples for testing. For MNIST, both the classical baseline SVM and the QSVM employ 64 features/qubits, whereas for Fashion MNIST, they use 96 features/qubits. Fig. 11 shows the confusion matrices for both datasets under SVM and QSVM, while Table 1 summarizes several macro-level performance metrics (accuracy, sensitivity, specificity, and $F_{1}$ score). For MNIST, the QSVM yields slightly higher accuracy, along with marginal improvements in sensitivity and specificity. A similar trend emerges for the more challenging Fashion-MNIST dataset, where QSVM also achieves higher overall accuracy and macro-level metrics than the classical SVM.
    These findings reinforce our earlier observations that, for moderate circuit sizes, QSVM can learn complex data distributions effectively. By embedding data points in a larger Hilbert space, the quantum kernel method can capture subtle features that improve class separability. However, as discussed previously, overfitting and the onset of Barren plateaus can degrade performance when the number of qubits becomes excessively large. Consequently, the design of circuit architectures and the choice of hyperparameters—particularly in multiclass settings—remain critical in balancing expressivity and generalization.
    Table 1: Macro Metrics for SVM and QSVM on MNIST and Fashion MNIST Datasets
    Dataset Model Accuracy Sensitivity Specificity F1-score
    MNIST SVM 0.8780 0.8820 0.9865 0.8783
    QSVM 0.8860 0.8900 0.9874 0.8864
    Fashion MNIST SVM 0.6140 0.6388 0.9578 0.6088
    QSVM 0.6660 0.6744 0.9633 0.6608
    4.3 Simulation with Single CPU and GPU
@@
    Figure 12: Benchmarking QSVM circuit simulation time using a single CPU and a single GPU. The GPU data shows minimal variation compared to the CPU’s scale.
    In this section, we compare the performance of a CPU and a GPU, as illustrated in Fig.12. To ensure a fair comparison, we employed Opt-Einsum for the contraction process on a single AMD EPYC 7J13 CPU, contrasting this with a single NVIDIA A100 GPU using cuTensorNet for the contraction process, with path reuse implemented. The detailed pseudocodes are discussed in Section 3.4. Moreover, it was necessary to synchronize the contraction paths in Opt-Einsum with those of cuTensorNet to ensure consistency. As depicted in Fig.12, the speedup provided by the GPU relative to the CPU becomes more pronounced as the number of simulated qubits increases. Consequently, for large-scale qubit simulations, GPUs demonstrate enhanced scalability and promise substantial benefits for future advanced qubit algorithms in simulation and emulation.
    5 Distributed simulation and Resource Estimation in HPC
@@
    In the final section of our study, a multi-GPU instance was utilized to expand the QSVM model via cuTensorNet to accommodate a dataset comprising 1,000 data points of 28x28 MNIST images. The implementation of multi-GPU resources to enhance quantum circuit simulation via cuStateVec is thoroughly detailed in the research conducted by Shaydulin et al. [41]. Our emphasis lies on leveraging the data from these experiments to rigorously assess both the computational costs and the temporal demands inherent in the tensor-network simulation of the QSVM algorithm within a multi-GPU processing framework.
    In our computational environment, each GPU within a node is interconnected using the high-bandwidth NVLink network and the NVIDIA Collective Communications Library (NCCL) to optimize intra-node communication. The input data is paired and evenly distributed across multiple GPUs via NCCL, where it is directly converted into a tensor network for computation. The results are then returned to a single GPU via NCCL to form the quantum kernel matrix for SVC classification. By harnessing these integrated technological benefits, we have successfully actualized the accelerated computational outcomes for managing large-scale qubit systems and complex datasets, as illustrated in Fig.13. Comparative analysis indicates that our performance metrics are on par with distributed simulation results documented in the existing scientific corpus, as cited in Bayraktar et. al.’s and Lykov et. al.’s work[24, 22].
    5.1 Benchmarking cuTensorNet Multi-GPU with MPI
    Fig. 13 illustrates the execution time required for quantum simulations in relation to the number of qubits. The data compares the performance of a single A100 GPU to systems utilizing 2, 4, and 8 GPUs in conjunction with MPI and within a single NVIDIA DGX node. It is evident from the results that the incorporation of multi-GPUs significantly decreases computation time, highlighting the strong linear speedup of cuTenserNet with MPI. The trend indicates a substantial reduction in execution time as the number of GPUs is increased, affirming the efficacy of multi-GPU setups in handling large datasets.
    5.2 Large Dataset Processing with Multi-GPU
@@
    6 Conclusion
    This paper has presented a comprehensive investigation into the feasibility and performance of large-scale QSVM simulations using a tensor-network-based framework integrated with NVIDIA’s cuQuantum SDK. By leveraging the cuTensorNet library on multi-GPU platforms, we significantly reduced the otherwise prohibitive computational overhead associated with simulating large qubit systems. Rigorous performance benchmarks demonstrated not only near-quadratic scaling for circuit simulations—thereby overcoming the exponential barriers of conventional state-vector approaches—but also robust speedups via MPI-based parallelization for quantum circuit simulation. Moreover, our experiments on benchmark datasets, including MNIST and Fashion-MNIST, revealed that QSVMs can achieve high classification accuracy, emphasizing the promise of quantum methods for complex, high-dimensional data. Crucially, the observed improvements in accuracy with increasing dataset size underscore the value of scalable simulation environments as a test bed for algorithmic refinements and real-world applications. The successful integration of cuTensorNet and multi-GPU infrastructures thus serves as a critical validation of quantum–HPC synergy, pointing to a practical route toward bridging near-term quantum hardware limitations and large-scale quantum machine learning goals. These results lay a foundation for further advances in high-performance quantum simulations and reinforce the potential impact of quantum-enhanced algorithms within the rapidly evolving Quantum–HPC ecosystem.
    Data Availability
    The code supporting the findings of this research is available on GitHub at the following repository: https://github.com/Tim-Li/cuTN-QSVM. This repository includes the scripts and data required to reproduce the results presented in this paper.
    Conflict of Interest Statement
    The authors declare that they have no conflict of interest related to this work.